Circular No. 8410

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

ASAS 002511+1217.2 A. Price, AAVSO, reports on the discovery of a new cataclysmic vari-

able (V = 10.49 ± 0.02) in Pisces by G. Pojmanski and the ASAS-3 survey on CCD images taken on Sept. 11.203 UT. Astrometry by A. Henden (1.0-m U.S. Naval Observatory telescope) gives its position as α = 0^{h}25^{m}11^{s}.087, δ = +12^{o}17′12′′.25 (equinox 2000.0; ± 0′′.04). The new variable is within the uncertainty boundaries of 1RXS J002510.8+121725 and 2MASS 00251111+1217121. Nothing is seen at that location on twentytwo ASAS-3 images taken between 2002 Dec. 15.041 and 2004 Sept. 6.363 (limiting V magnitude 14). The variable is detected at mag 17.5 on a Digitized Sky Survey blue scan of a plate taken on 1992 Aug. 29. Fourier (+ CLEANEST) analysis of CCD observations by L. Cook (Concord, CA; unfiltered), A. Oksanen (Muurame, Finland; V), D. Boyd (West Challow, Wantage, Oxfordshire, England; V), and M. Koppelman (Golden Valley, MN; R_{c}) between Sept. 15.902 and 19.300 reveal superhumps with a period of 0.05689 ± 0.00001 day. The superhump period and lack of previously detected outbursts suggests classification as a U-Gem/WZ-Sge-type dwarf nova. Additional magnitudes: Sept. 15.824, 10.7 (P. Schmeer, Bischmisheim, Germany; visual); 19.868, 11.4 (E. Muyllaert, Oostende, Belgium; visual); 20.459, 11.58 ± 0.01 (Cook; unfiltered CCD).

SUPERNOVA 2003jd K. S. Kawabata, Hiroshima University; K. Maeda, J. Deng and K.

Nomoto, University of Tokyo; P. A. Mazzali and E. Pian, Trieste Observatory; L. Wang, Lawrence Berkeley National Laboratory; and Y. Ohyama and M. Iye, National Astronomical Observatory of Japan, write on behalf of the Subaru supernova ToO group: “We obtained an optical spectrum (range 380–720 nm) of the peculiar type-Ic supernova 2003jd (IAUC 8232, 8234) on Sept. 12.4 UT with the Subaru 8.2-m telescope (+ FOCAS). The spectrum still shows an overall similarity to that of the hypernova 1998bw (at 201 days; Patat et al. 2001, Ap.J. 555, 900). However, the broad [O I] 630.0- and 636.3-nm emission feature seems to have a double-peaked profile with a separation of ∼ 10 nm. This type of profile is typical of a disk-like or torus structure viewed near the equatorial plane. The broad emission line at 450–460 nm, possibly Mg I] 457.1 nm, barely shows a similar profile. The observed spectrum is contaminated by strong narrow emission lines and a blue continuum excess, which may be ascribed to an H II region in the host galaxy and/or a background star-forming galaxy.”

2004 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green